# Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention

### v3

@@ 8 Discussion
Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and human spontaneous thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI. In an era of growing concern about AI’s cognitive resemblance to humans, our framework provides preliminary results differentiating these phenomena, thereby opening pathways for future interdisciplinary research at the interface of artificial and human cognition.
Limitations. Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from Li et al., 2024b and defining topics as TPGs. These abstractions do not fully capture the complexities of contemporary LLMs, including deep attention architectures, alternative loss functions, and diverse training objectives. Despite loosening these assumptions, our experiments suggest that the essence of our core theoretical conclusions holds across modern LLMs within our framework of study. Future work will investigate how broadly these theoretical insights generalize to complex LLM architectures.
@@ A.2 Proof of Lemma 6
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$ or $i=j$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@
Since $\bf X$ belongs to $\mathbb{T}$, $\forall{\bf x}\in{\bf X}$ we have $x\in\mathcal{G}^{(\bar{x})}$, therefore from the construction of TPGs by Li et al., 2024b , for every ${\bf x}_{i},{\bf x}_{j}\in{\bf X}$ we have one of the these relationships: $(x_{i}\Rightarrow x_{j})$, $(x_{j}\Rightarrow x_{i})$, $(x_{i}\asymp x_{j})$ or $x_{i}=x_{j}$. From the constraints in Algo-GD:
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}=0\iff(x_{j}\asymp x_{i})\in\mathcal{G}^{(\bar{x})}\text{ or }i=j$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}>1\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@ A.4 Proof of Theorem 3
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ or $i=j$, then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $i=j\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
@@ A.6 Proof of Theorem 5
    Let ${\bf X}_{l}^{\prime}=[{\bf x}_{1,l}^{\prime}\ {\bf x}_{2,l}^{\prime}\ \cdots\ {\bf x}_{T,l}^{\prime}]^{\top}$ be a random sequences generated as described in the theorem, where the size of $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ is $l$. Let $k_{i,l}$ be the number of times ${\bf x}_{i}$ is selected in ${\bf X}_{l}^{\prime}$ Let $A_{l}=\max_{1\leq i\leq l}k_{i,l}$ and $B_{l}=\max_{l+1\leq i\leq K}k_{i,l}$. Let $P(l)=\mathbb{P}(B_{l}>A_{l})$. We want to prove $P(l+1)\leq P(l)$. We construct a coupling between ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ by performing $T$ independent trials. For each trial $i$ we generate a uniform random variable $U_{i}$ in $[0,1]$ and we choose tokens in ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ in this way:
@@
Appendix B Detailed Explanation of Assumption 4
@@
Appendix C Explanation of TPGs Combination
Figure 3 presents $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. In $\mathcal{G}^{(4)}_{ab}$ for the mixed-topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$. The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
Appendix D Detailed Simulation Studies with Single-layer Self-attention
D.1 Simulation Process
@@
Trained attention weights. We employ a single-layer attention mechanism implemented in PyTorch. The model is trained using the SGD optimizer with a learning rate $\eta=0.01$ for $8000$ iterations. The training of attention weights is divided into two stages for each instance: (1) computing ${\bf W}^{\text{svm}}$ for each topic  [^4] Note: ${\bf W}^{\text{svm}}={\bf 0}$ means the number of SCCs is $1$ for $\mathcal{G}^{k},\forall k\in[K]$. During the simulation, we proceed to the next instance when ${\bf W}^{\text{svm}}={\bf 0}$ until reaching a total of $100$ instances.; (2) get $\bf{W}(\tau)$ at each iteration for each topic. In Stage (1), prior to using the CVXPY package to get ${\bf W}^{\text{svm}}$, SCCs are identified for each TPG derived from the using Tarjan’s algorithm. Afterward, ${\bf W}^{\text{svm}}$ is normalized to ensure consistency in subsequence computations. In Stage (2), the $MLayerAttn$ function encapsulates the architecture of a single-layer attention-based model. The training function is then used to optimize the attention weights by minimizing the loss defined in ERM. Finally, the correlation between ${\bf W}^{\text{svm}}$ and $\bf{W}(\tau)$ is calculated using the dot product.
Next token prediction. To differentiate the input sequence length of the testing data from that of the training data, we introduce $T_{test}$. TPGs based on the training dataset DSETa are utilized to generate test datasets consisting of 100 sequences from $\mathbb{T}_{a}$ per epoch. Specifically, the last token ${\bf x}_{T_{test}}$ of the test input sequence is randomly selected from $K$ tokens (i.e. ${\bf x}_{T_{test}}={\bf e}_{k}$) and the remaining input tokens are randomly chosen based on the SCCs of $\mathcal{G}^{k}_{a}$, where tokens with higher priority are assigned greater weights. For instance, in $\mathcal{G}^{4}_{a}$, tokens ${\bf e}_{1},{\bf e}_{2},{\bf e}_{4}$ are captured with the priority order ${\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. The weights assigned to input tokens ${\bf e}_{1},{\bf e}_{2}$, and ${\bf e}_{4}$ are $0.4,0.4$, and $0.2$, respectively. It reflects that ${\bf e}_{1}$ and ${\bf e}_{2}$ are in the same higher-priority SCC, thus having greater weights compared to ${\bf e}_{4}$. Intuitively, tokens within the same SCC are more likely to co-occur than those from different SCCs. This approach enables the generated test input sequences to mimic real word relationships and reflect their contextual groupings. Following the generation of the test dataset from $\mathbb{T}_{a}$, the next tokens $\hat{y}_{\bf w_{a}}$ and $\hat{y}_{\bf w_{ab}}$ are predicted by Equation 1, with ${\bf W}_{a}$ and ${\bf W}_{ab}$ obtained from the last iteration. To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs. Our code for all simulations is available in ( ).
D.2 Additional Experiments to Support Theorem 3
@@
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 7. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (simulation details in Appendix D). Figure 7 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.

### v4

@@ 8 Discussion
Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and spontaneous human thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI. In an era of growing concern about AI’s cognitive resemblance to humans, our framework provides preliminary results differentiating these phenomena, thereby opening pathways for future interdisciplinary research at the interface of artificial and human cognition.
Limitations.
Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from li2024mechanics and defines topics as TPGs. These abstractions do not fully capture the complexities of contemporary LLMs, including deep attention architectures, alternative loss functions, and diverse training objectives. Despite loosening these assumptions, our experiments suggest that the essence of our core theoretical conclusions holds across modern LLMs within our framework of study. Future work will investigate how broadly these theoretical insights generalize to complex architectures, for example within longer context windows and/or LLM outputs.
Code.
Acknowledgments and Disclosure of Funding
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant DGECR-2022-04531. The authors thank Prof. Yingcong Li, the anonymous reviewers, and the session chair for their valuable feedback and insightful suggestions, which greatly improved the quality of this work.
Supplementary Materials
@@ A.2 Proof of Lemma 2
  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$ or $i=j$.
  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@
Since $\bf X$ belongs to $\mathbb{T}$, $\forall{\bf x}\in{\bf X}$ we have $x\in\mathcal{G}^{(\bar{x})}$, therefore from the construction of TPGs by li2024mechanics, for every ${\bf x}_{i},{\bf x}_{j}\in{\bf X}$ we have one of the these relationships: $(x_{i}\Rightarrow x_{j})$, $(x_{j}\Rightarrow x_{i})$, $(x_{i}\asymp x_{j})$ or $x_{i}=x_{j}$. From the constraints in Algo-GD:
[…]
  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}>1\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@ A.4 Proof of Theorem 2
  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ or $i=j$, then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $i=j\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
@@ A.6 Proof of Theorem 4
    Let ${\bf X}_{l}^{\prime}=[{\bf x}_{1,l}^{\prime}\ {\bf x}_{2,l}^{\prime}\ \cdots\ {\bf x}_{T,l}^{\prime}]^{\top}$ be a random sequences generated as described in the theorem, where the size of $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ is $l$. Let $k_{i,l}$ be the number of times ${\bf x}_{i}$ is selected in ${\bf X}_{l}^{\prime}$ . Let $A_{l}=\max_{1\leq i\leq l}k_{i,l}$ and $B_{l}=\max_{l+1\leq i\leq K}k_{i,l}$. Let $P(l)=\mathbb{P}(B_{l}>A_{l})$. We want to prove $P(l+1)\leq P(l)$. We construct a coupling between ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ by performing $T$ independent trials. For each trial $i$ we generate a uniform random variable $U_{i}$ in $[0,1]$ and we choose tokens in ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ in this way:
@@
Appendix B Explanation of Assumption 4
@@
Appendix C Detailed simulation studies with single-layer self-attention
C.1 Simulation process
@@
Trained attention weights. We employ a single-layer attention mechanism implemented in PyTorch. The model is trained using the SGD optimizer with a learning rate $\eta=0.01$ for $8000$ iterations. The training of attention weights is divided into two stages for each instance: (1) computing ${\bf W}^{\text{svm}}$ for each topic; [^4] Note: ${\bf W}^{\text{svm}}={\bf 0}$ means the number of SCCs is $1$ for $\mathcal{G}^{k},\forall k\in[K]$. During the simulation, we proceed to the next instance when ${\bf W}^{\text{svm}}={\bf 0}$ until reaching a total of $100$ instances. (2) get $\bf{W}(\tau)$ at each iteration for each topic. In Stage (1), prior to using the CVXPY package to get ${\bf W}^{\text{svm}}$, SCCs are identified for each TPG derived from the using Tarjan’s algorithm. Afterward, ${\bf W}^{\text{svm}}$ is normalized to ensure consistency in subsequence computations. In Stage (2), the $MLayerAttn$ function encapsulates the architecture of a single-layer attention-based model. The training function is then used to optimize the attention weights by minimizing the loss defined in ERM. Finally, the correlation between ${\bf W}^{\text{svm}}$ and $\bf{W}(\tau)$ is calculated using the dot product.
Next token prediction. To differentiate the input sequence length of the testing data from that of the training data, we introduce $T_{test}$. TPGs based on the training dataset DSETa are utilized to generate test datasets consisting of 100 sequences from $\mathbb{T}_{a}$ per epoch. Specifically, the last token ${\bf x}_{T_{test}}$ of the test input sequence is randomly selected from $K$ tokens (i.e. ${\bf x}_{T_{test}}={\bf e}_{k}$) and the remaining input tokens are randomly chosen based on the SCCs of $\mathcal{G}^{k}_{a}$, where tokens with higher priority are assigned greater weights. For instance, in $\mathcal{G}^{4}_{a}$, tokens ${\bf e}_{1},{\bf e}_{2},{\bf e}_{4}$ are captured with the priority order ${\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. The weights assigned to input tokens ${\bf e}_{1},{\bf e}_{2}$, and ${\bf e}_{4}$ are $0.4,0.4$, and $0.2$, respectively. It reflects that ${\bf e}_{1}$ and ${\bf e}_{2}$ are in the same higher-priority SCC, thus having greater weights compared to ${\bf e}_{4}$. Intuitively, tokens within the same SCC are more likely to co-occur than those from different SCCs. This approach enables the generated test input sequences to mimic real word relationships and reflect their contextual groupings. Following the generation of the test dataset from $\mathbb{T}_{a}$, the next tokens $\hat{y}_{\bf w_{a}}$ and $\hat{y}_{\bf w_{ab}}$ are predicted by Equation 1, with ${\bf W}_{a}$ and ${\bf W}_{ab}$ obtained from the last iteration. To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs.
C.2 Additional experiments to support Theorem 2
@@
The attention priority similarity quantifies how well the weights $\bf W$$^{\prime}$ preserve the attention priority of the weights $\bf W$}. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 7. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as li2024mechanics. We generate test sequences $\bf Z$} within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C). Figure 7 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 2.